OpenStudy (anonymous):

in scientific notation , what is the sum of 2.3 X 10-3 and 5.4 x 10 -4

OpenStudy (aum):

\( \Large 2.3 \times 10^{-3} ~~+~~ 5.4 \times 10^{-4} = \\ \Large 2.3 \times \frac{10}{10} \times 10^{-3} ~~+~~ 5.4 \times 10^{-4} = \\ \Large 23 \times 10^{-4} ~~+~~ 5.4 \times 10^{-4} = 28.4 \times 10^{-4} \)